# The Internal Model Principle of Time-Varying Optimization

### v1

@@ VII-B Application to solve the dynamic traffic assignment problem in transportation
$\quad+\sum_{v\in\mathcal{V}}\lambda_{v}\left(\delta_{v}(\theta(t))-\sum_{\begin{subarray}{c}j\in\mathcal{E}:\\ j^{-}=v\end{subarray}}x_{j}+\sum_{\begin{subarray}{c}j\in\mathcal{E}:\\ j^{+}=v\end{subarray}}x_{j}\right),$
where $\lambda:=(\lambda_{1},\dots,\lambda_{|\mathcal{V}|})$ is the vector of Lagrange multipliers. We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section VI); the inequality constraints in (VII-B) have been accounted for by projecting $x(t)$ onto the feasible set. Here, matrix $L$ has been chosen so that the eigenvalues of $S-LQ$ are uniformly distributed in the open real interval $(-1,-2)$. Notice that, since the latencies are strictly increasing, the Lagrangian is strongly convex-strongly concave [37], and thus the problem admits a unique critical point that is a saddle point. It follows that our algorithm is guaranteed to converge to a minimizer of (VII-B). Simulation for this problem are presented in Fig. 5. The simulation shows that $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty$ (see Fig. 5-second figure) and $y(t)\rightarrow 0,$ which implies that the algorithm successfully computes the critical point (see Fig. 5-third figure). Finally, the bottom figure of Fig. 5 illustrates the rate of convergence of the algorithm, which, as expected, is governed by the placement of the eigenvalues of the observer.
Figure 5: Simulation results illustrating the performance of the optimization algorithm synthesized using Algorithm 1 to solve the dynamic traffic assignment problem (VII-B). See Section VII-B for a discussion. The bottom figure show that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium. Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of $S-LQ.$
VIII Conclusions
We showed that the problem of designing optimization algorithms for time-varying optimization problems can be reconducted to the output regulation problem of nonlinear multivariable systems. This connection allowed us to prove the internal model principle of time-varying optimization, which states that asymptotic tracking can be achieved only if the optimization algorithm incorporates a reduplicated model of the temporal variability of the problem. Further, we showed that asymptotic tracking can be achieved under more relaxed assumptions (namely, assumption (A2)) than what is normally imposed in the literature (see assumption (A3)). Moreover, our algorithm structure is novel in the literature, and it relies on the use of an observer for the temporal variability of the problem. This work opens the opportunity for several directions of future work, including an investigation at discrete-time and possibly the application in feedback optimization.
@@ VIII Conclusions
where $f$ is a $C^{k}$ vector field defined on an open subset $U$ of n, and let $x_{\circ}\in U$ be an equilibrium point for $f$, i.e., $f(x_{\circ})=0$. Without loss of generality, suppose $x^{\circ}=0$. Let $F=\left[\frac{\partial f}{\partial x}\right]_{x=0},$ denote the Jacobian matrix of $f$ at $x=0$. Suppose the matrix $F$ has $n^{\circ}$ eigenvalues with zero real part, $n^{-}$ eigenvalues with negative real part, and $n^{+}$ eigenvalues with positive real part. Let $E^{-},E^{\circ},$ and $E^{+}$ be the (generalized) real eigenspaces of $F$ associated with eigenvalues of $F$ lying on the open left half plane, the imaginary axis, and the open right half plane, respectively. Note that $E^{\circ},E^{-},E^{+}$ have dimension $n^{\circ},n^{-},n^{+}$, respectively and that each of these spaces is invariant under the flow of $\dot{x}=Fx.$ If the linear mapping $F$ is viewed as a representation of the differential (at $x=0$) of the nonlinear mapping $f$, its domain is the tangent space $T_{0}U$ to $U$ at $x=0$, and the three subspaces in question can be viewed as subspaces of $T_{0}U$ satisfying $T_{0}U=E^{\circ}\oplus E^{-}\oplus E^{+}.$ We refer to [39, Sec. A.II] for a precise definition of $C^{k}$ manifolds; loosely speaking, a set $S\subset U$ is a $C^{k}$ manifold it can be locally represented as the graph of a $C^{k}$ function.
@@
A $C^{k}$ manifold $S$ of $U$ is said to be locally invariant for (26) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (26) satisfying $x(0)=x_{\circ}$ is such that $x(t)\in S$ for all $t\in\left(t_{1},t_{2}\right)$. $\square$
@@
a curve $y=\pi(\theta)$ is an invariant manifold for (27) if the solution of (27) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0$ The notion of invariant manifold is useful as, under certain assumptions, it allow us to reduce the analysis of (26) to the study of a reduced system in the variable $\theta$ only. The remainder of this section is devoted to formalizing this fact.
@@
Returning to the decomposition (27), intuitively, the invariant manifold $y=\pi(\theta)$ is said to be a center manifold when all orbits of $y$ decay to zero and those of $\theta$ neither decay nor grow exponentially.
In what follows, we will assume that all eigenvalues of $F$ have nonpositive real part, i.e.,
$n^{+}=0.$ (28)
When (28) holds, it is always possible choose coordinates in $U$ such that (26) reads as:
$\dot{y}=Ay+g(y,\theta)$ (29a)
$\dot{\theta}=B\theta+h(y,\theta)$ (29b)
@@
Assume that (28) holds. There exist a neighborhood $V\subset{}^{n^{0}}$ of 00 and a class $C^{k-1}$ mapping $\pi:V\rightarrow{}^{n^{-}}$ such that the set
@@
as deduced by differentiating with respect to time any solution $(y(t),\theta(t))$ of (29) on the manifold $y(t)=\pi(\theta(t))$. In other words, any center manifold for (29) can equivalently be described as the graph of a mapping $y=\pi(\theta)$ satisfying the partial differential equation (31), with the constraints specified by (30).
@@
Theorem 8 shows existence but not uniqueness of a center manifold. Moreover, (i) if $g$ and $h$ are $C^{k},k\in\mathbb{N}_{>0},$ then (29) admits a $C^{k-1}$ center manifold; (ii) if $g$ and $h$ are $C^{\infty}$ functions, then (29) has a $C^{k}$ center manifold for any finite $k$, but not necessarily a $C^{\infty}$ center manifold. $\square$
The next result shows that any $y$-trajectory of (29), starting sufficiently close to the origin, converges, as time tends to infinity, to a trajectory that belongs to the center manifold.
@@
Assume that (28) holds and suppose $y=\pi(\theta)$ is a center manifold for (29) at $(0,0)$. Let $(y(t),\theta(t))$ be a solution of (29). There exists a neighborhood $U^{\circ}$ of $(0,0)$ and real numbers $M>0$ and $K>0$ such that, if $(y(0),\theta(0))\in U^{\circ}$, then for all $t\geq 0$,
@@
From the above discussion, any trajectory of (29) starting at a point $y^{\circ}=\pi\left(\theta^{\circ}\right)$ of a center manifold satisfies:
$y(t)=\pi(\zeta(t)),$ $\theta(t)=\zeta(t),$
where $\zeta(t)$ is any solution of
$\dot{\zeta}=B\zeta+h(\pi(\zeta),\zeta),$ (32)
satisfying the initial condition $\zeta(0)=\theta_{\circ}$. This decomposition allows us to predict the asymptotic behavior of (29) by studying the asymptotic behavior of a reduced-order system, namely, (32). This is formalized in the following result.

### v2

@@ 7.3 Application to solve the dynamic traffic assignment problem in transportation
where $\lambda:=(\lambda_{1},\dots,\lambda_{|\mathcal{V}|})$ is the vector of Lagrange multipliers. We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section 6); the inequality constraints in (7.3) have been accounted for by projecting $x(t)$ onto the feasible set. Here, matrix $L$ has been chosen so that the eigenvalues of $S-LQ$ are uniformly distributed in the open real interval $(-1,-2)$. Notice that, since the latencies are strictly increasing, the Lagrangian is strongly convex-strongly concave [58], and thus the problem admits a unique critical point that is a saddle point. It follows that our algorithm is guaranteed to converge to a minimizer of (7.3). Simulations for this problem are presented in Fig. 6. The simulation shows that $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty$ (see Fig. 6-second figure) and $y(t)\rightarrow 0$ (up to machine precision) which implies that the algorithm successfully computes a critical point (see Fig. 6-third figure). Finally, the bottom figure of Fig. 6 illustrates the rate of convergence of the algorithm, which, as expected, is governed by the placement of the eigenvalues of the observer.
8 Conclusions
We showed that the problem of designing algorithms for time-varying optimization problems can be reformulated as a nonlinear multivariate output regulation problem. This connection allowed us to prove the internal model principle of time-varying optimization, which states that (when the algorithm has access only to functional evaluations of the gradient) asymptotic tracking can be achieved only if the algorithm incorporates a reduplicated model of the temporal variability of the problem. On the other hand, when the time-varying parameters are measurable, asymptotic tracking can be achieved by seeking a mapping that zero the gradient. Importantly, our results show that asymptotic tracking can be achieved under more relaxed assumptions than what is normally imposed in the literature. Moreover, our algorithm structure is novel in the literature, and it relies on the use of an observer for the temporal variability of the problem. This work opens the opportunity for several directions of future work, including the relaxation of the convexity and smoothness assumptions, an investigation of discrete-time problems, and applications to feedback optimization.
@@ 8 Conclusions
where $f$ is a $C^{k}$ vector field defined on an open subset $U$ of n, and let $x_{\circ}\in U$ be an equilibrium point for $f$, i.e., $f(x_{\circ})=0$. Without loss of generality, suppose $x^{\circ}=0$. Let $F=\left[\frac{\partial f}{\partial x}\right]_{x=0},$ denote the Jacobian matrix of $f$ at $x=0$. Suppose the matrix $F$ has $n^{\circ}$ eigenvalues with zero real part, $n^{-}$ eigenvalues with negative real part, and $n^{+}$ eigenvalues with positive real part. Let $E^{-},E^{\circ},$ and $E^{+}$ be the (generalized) real eigenspaces of $F$ associated with eigenvalues of $F$ lying on the open left half plane, the imaginary axis, and the open right half plane, respectively. Note that $E^{\circ},E^{-},E^{+}$ have dimension $n^{\circ},n^{-},n^{+}$, respectively and that each of these spaces is invariant under the flow of $\dot{x}=Fx.$ If the linear mapping $F$ is viewed as a representation of the differential (at $x=0$) of the nonlinear mapping $f$, its domain is the tangent space $T_{0}U$ to $U$ at $x=0$, and the three subspaces in question can be viewed as subspaces of $T_{0}U$ satisfying $T_{0}U=E^{\circ}\oplus E^{-}\oplus E^{+}.$ We refer to [60, Sec. A.II] for a precise definition of $C^{k}$ manifolds; loosely speaking, a set $S\subset U$ is a $C^{k}$ manifold it can be locally represented as the graph of a $C^{k}$ function.
@@
A $C^{k}$ manifold $S$ of $U$ is locally invariant for (37) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (37) satisfying $x(0)=x_{\circ}$ satisfiest $x(t)\in S$ for all $t\in(t_{1},t_{2})$. $\square$
@@
a curve $y=\pi(\theta)$ is an invariant manifold for (38) if the solution of (38) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0$ The notion of invariant manifold is useful as, under certain assumptions, it allows us to reduce the analysis of (37) to the study of a reduced system in the variable $\theta$ only. The remainder of this section is devoted to formalizing this fact.
@@
Intuitively, the invariant manifold $y=\pi(\theta)$ is a center manifold for (38) when all orbits of $y$ decay to zero and those of $\theta$ neither decay nor grow exponentially.
In what follows, we will assume that all eigenvalues of $F$ have nonpositive real part, i.e., $n^{+}=0$. When this holds, it is always possible to choose coordinates in $U$ such that (37) is
$\dot{y}=Ay+g(y,\theta),$ (39a)
$\dot{\theta}=B\theta+h(y,\theta),$ (39b)
@@
Assume that $n^{+}=0$. There exists a neighborhood $V\subset{}^{n^{0}}$ of 00 and a class $C^{k-1}$ mapping $\pi:V\rightarrow{}^{n^{-}}$ such that the set
@@
as deduced by differentiating with respect to time any solution $(y(t),\theta(t))$ of (39) on the manifold $y(t)=\pi(\theta(t))$. In other words, any center manifold for (39) can equivalently be described as the graph of a mapping $y=\pi(\theta)$ satisfying the partial differential equation (41), with the constraints (40).
@@
Theorem 9 shows existence but not the uniqueness of a center manifold. Moreover, (i) if $g$ and $h$ are $C^{k},k\in\mathbb{N}_{>0},$ then (39) admits a $C^{k-1}$ center manifold; (ii) if $g$ and $h$ are $C^{\infty}$ functions, then (39) has a $C^{k}$ center manifold for any finite $k$, but not necessarily a $C^{\infty}$ center manifold. $\square$
The next result shows that any $y$-trajectory of (39), starting sufficiently close to the origin converges, as time tends to infinity, to a trajectory that belongs to the center manifold.
@@
Assume that $n^{+}=0$ and suppose $y=\pi(\theta)$ is a center manifold for (39) at $(0,0)$. Let $(y(t),\theta(t))$ be a solution of (39). There exists a neighborhood $U^{\circ}$ of $(0,0)$ and real numbers $M>0$ and $K>0$ such that, if $(y(0),\theta(0))\in U^{\circ}$, then for all $t\geq 0$,
@@
From the above discussion, any trajectory of (39) starting at a point $y^{\circ}=\pi\left(\theta^{\circ}\right)$ of a center manifold satisfies $y(t)=\pi(\zeta(t))$ and $\theta(t)=\zeta(t)$, where $\zeta(t)$ is any solution of
$\dot{\zeta}=B\zeta+h(\pi(\zeta),\zeta),\qquad\zeta(0)=\theta_{\circ}.$ (42)
This decomposition allows us to predict the asymptotic behavior of (39) by studying the asymptotic behavior of the reduced-order system (42). This is formalized in the following result.